A sound wave has a frequency of 658 Hz in air
and a wavelength of 0.53 m.
What is the temperature of the air? Assume the velocity of sound at 0◦C is 335 m/s.
Answer in units of ◦C.
1
Expert's answer
2020-02-18T09:36:30-0500
"v=\lambda f=(0.53)(658)=348.7\frac{m}{s}"

"\frac{T}{T_0}=\frac{v^2}{v_0^2}-1"

"\frac{T}{273}=\frac{348.7^2}{335^2}-1"

"T=22.8\degree C"